# Autonomous Exploration and Semantic Updating of Large-Scale Indoor Environments with Mobile Robots

### v2

@@ IV Experiments
We have conducted experiments in two different environments with different object categories and evaluate the performance of our system in those environments.
@@
Environment A. The fourth floor of the ECSS building at UT Dallas is a large indoor environment measuring approximately $93m\times 90m$. It consists of 18 corridors with a total traversal length of approximately $800m$. These corridors were mostly furnished with typical office furniture, including doors, tables, and chairs. Therefore, we only consider these 3 classes, i.e., table, door and chair for semantic mapping of this environment.
@@
Environment B. This environment is a laboratory environment with a reduced size measuring $9m\times 13m$. This space included a more diverse set of objects such as televisions, trash bins, umbrellas, chairs, cabinets and persons.
@@
Robot. We run the experiments on a Fetch mobile robot equipped with a LiDAR sensor and an RGB-D camera as depicted in Fig. 6.
Figure 6: The Fetch robot maps the environment using a $220^{o}$ laser scanner of $25m$ range, while leveraging depth observations for obstacle avoidance. Its RGB-D camera operates at 30fps, enabling real-time object detection and segmentation for semantic mapping running on the connected GPU laptop.
@@
Exploration and Mapping. To autonomously explore and map the environments, we integrate the mapping, exploration, and navigation modules. The Gmapping ROS package [37] was used to incrementally build a 0.1m/pixel occupancy grid map, using laser scan data from the Fetch robot. Simultaneously, we ran our custom-built Dynamic Window Frontier Exploration module to identify and select best valued frontier in the search window. The move_base ROS package [40] guided the robot to these frontiers by combining a global planner for long term path planning and a local planner that dynamically avoids obstacles not seen during global planning. This ensures safe navigation when any new objects are added or moved around in the environment. The robot’s maximum velocity was limited to 0.6m/s, with a 0.7m inflation radius to prevent it from traveling too close to the obstacles. In Environment A, after 150 minutes of exploration, which also included several stops due to close proximity of persons in narrow corridors, the environment was fully mapped, and 2,250 trajectory points were recorded at 0.25Hz. In Environment B, exploration was completed in approximately 4 minutes, with 212 points recorded at 1Hz.
Environment Traversing. The greedy TSP described in Sect. III-B1 is formulated using the poses recorded during the exploration, resulting in traversal trajectories of 130 poses for Environment A and 15 poses for Environment B. In particular, the time to revisit is significantly reduced compared to the exploration phase. Revisiting took approximately 35 minutes for Environment A and 2 minutes for Environment B. This highlights the efficiency gained through this traversal optimization.
Real-time Construction and Updating of Semantic Maps. To enable real-time semantic mapping, the Fetch robot was equipped with a laptop featuring NVIDIA RTX 4090 GPU, running Ubuntu 20.04 and ROS Noetic. The laptop and Fetch robot were connected via Ethernet for subscribing and publishing the data through ROS topics.
Initially, the environment remained unchanged compared to its state during exploration. The system then begins with the built occupancy map. After localizing itself using AMCL ROS [40], the robot progressively builds the semantic map. Subsequently, semantic map updates were tested by removal, addition, and/or moving the objects in the environments. Visual results of environment A can be found in Fig. 2. More details of it can be found in the supplementary video. Fig. 7 depicts some of the changes in Environment B, reflected in the semantic updating stages.
@@
Figure 7: Visualization of some selected environmental changes in Environment B across two update stages. In Update stage I, 2 trash bins are removed. After that in update stage II, 2 chairs were removed. When robot revisited, these changes are reflected in the updated semantic map.
Env Class Construction Update I Update II
GT Map GT Map GT Map
A Table 45 45 45 39 45* 37
Chair 109 84 99 73 109* 69
Door 170 153 170 155 170 156
B Person 2 2 2* 2 2 2
Trash Bin 2 2 0 0 0 1
Umbrella 1 1 0 0 0 0
Chair 3 3 3 3 1* 1
Television 1 2 1 1 1 1
Cabinet 3 2 3 1 3 2
TABLE II: Semantic map object count across the construction and update stages for A & B Environments. GT: Ground Truth, Map: no. of objects mapped. The table compares the ground truth with the actual results across the construction, update I, and update II scenarios. * indicates that some objects have been relocated from their earlier position.
IV-A Quantitative Analysis
Table II summarizes the number of objects detected in both Environment A and Environment B during the Construction, Update I and Update II phases. In Environment A, we can see that while doors and tables were stably detected, the detection of chairs was relatively more challenging. During update I, the removal of 10 chairs after the construction phase is reflected in the decrease in the number of chairs detected from 84 to 73. However, in Update II, only 69 chairs were detected, with false negatives increasing to 29 as shown in Table III. This was mainly because some relocated chairs were placed in the path of the robot traversal trajectory. As a result, robot took a detour to avoid hitting them, causing it to miss some chairs that were not relocated and previously detected. Moreover, people occupying chairs occluded them, increasing false negatives. In contrast, Environment B has very stable detections, with many classes having 100% precision. However, some categories like cabinets that are dark colored, are challenging to detect in nominal lightning, resulting in them not being detected.
IV-B Qualitative Analysis
@@
Class Construction Update I Update II
TP FP FN Precision TP FP FN Precision TP FP FN Precision
Environment A
Table 36 9 4 0.9 36 3 9 0.8 33 4 6 0.85
Chair 84 0 16 0.84 72 1 23 0.76 67 2 29 0.69
Door 140 13 28 0.83 140 15 26 0.84 142 14 28 0.84
Environment B
Person 2 0 0 1 2 0 0 1 2 0 0 1
Trash Bin 2 0 0 1 0 0 0 - 0 1 0 0
Umbrella 1 0 0 1 0 0 0 - 0 0 0 -
Chair 3 0 0 1 3 0 0 1 1 0 0 1
Television 1 1 0 0.5 1 0 0 1 1 0 0 1
Cabinet 2 0 1 0.67 1 0 2 0.33 2 0 1 0.67
TABLE III: Semantic map object count across the construction and update stages for A & B Environments. TP: True Positives, FP: False Positives, FN: False Negatives, Precision: Precision for detected objects across the construction and update stages.
@@
Limitations in Constrained Spaces. Narrow corridors limit the robot’s view of objects like doors, as seen in Fig. 9(f). Several of such doors are not detected, leading to a reduction in the number of object nodes in the semantic map. This is evident in Table II.
@@ V Conclusion and Future Work
We introduce a robotic system that enables a mobile robot to autonomously explore an unknown environment, recognize objects in the environment, and subsequently build a semantic map of the environment. The system utilizes a hybrid representation of the semantic map that consists of an occupancy grid map for geometry and a topological map for semantics. This representation enables us to update the semantic map efficiently to reflect the changes in the environment. Experiments conducted on a mobile manipulatory system in a large- and medium-scale indoor environments demonstrate the effectiveness of our system in autonomous exploration and semantic mapping.
@@ Acknowledgment
This work was supported by the DARPA Perceptually-enabled Task Guidance (PTG) Program under contract number HR00112220005 and the Sony Research Award Program. The work of T. Summers was supported by the United States Air Force Office of Scientific Research under Grant FA9550-23-1-0424 and the National Science Foundation under Grant ECCS-2047040. We would like to thank our colleague, Jishnu Jaykumar P, for his assistance during the experiments.

### v3

@@ IV Experiments
To evaluate the effectiveness of our proposed Dynamic Window Frontier Exploration method (Alg. 1), we conducted experiments in a large-scale simulated hospital environment[41] of approximately $1440~\text{sq.m}$, in the Gazebo simulator. We conducted 20 iterations of exploration, where in each iteration a Fetch mobile manipulator robot was spawned at a different starting location and explored for 15 minutes. The same starting locations are used for both our approach and conventional frontier exploration. During each run, we recorded the area covered and the path length traveled, allowing a direct comparison between our approach and the conventional frontier exploration baseline [37].
[…]
Baseline[37] 669.12 107.10
DFWE (ours) 721.05 94.30
TABLE II: Average area coverage and path length for baseline vs. DFWE in a large-scale simulated environment.
The results in Table II demonstrate the improved exploration efficiency using our method. Compared to the baseline exploration, our method explores $\mathbf{\approx 7.8\%}$ more area while traveling $\mathbf{\approx 12\%}$ less on average. These numbers are expected to get better as the size of environment increases.
@@
We conducted semantic mapping experiments in two different real-world environments with different object categories and evaluated the performance of our system.
@@
Environment A. This is a large indoor environment measuring approximately $\mathbf{93m\times 90m}$. It consists of 18 corridors with a total traversal length of approximately 800 m These corridors were mostly furnished with typical office furniture, including doors, tables, and chairs. Therefore, we only consider these 3 classes, i.e., table, door and chair for reliable semantic mapping of this environment.
@@
Environment B. This is a laboratory environment with a reduced size measuring $9m\times 13m$. This space included a more diverse set of objects such as televisions, trash bins, umbrellas, chairs, cabinets and persons. Only objects with a detection confidence of 0.6 or higher are considered.
@@
Robot. We run the experiments on a Fetch mobile robot equipped with a LiDAR sensor and an RGB-D camera.
IV-A Real World Exploration and Mapping
@@
We integrate mapping, exploration, and navigation to autonomously map environments. Using Gmapping [38], a 0.1 m/pixel resolution occupancy map was built using laser scans, while our Dynamic Window Frontier Exploration module selected the best frontiers in the search radius. The move_base ROS package [42] guided the robot safely to these frontiers with global and local planning, dynamically avoiding unexpected obstacles. The robot was limited to 0.6 m/s speed with a 0.7 m inflation radius for safe navigation in the environment. In Environment A, 150 min of the exploration (including stops for people in narrow corridors) produced a full map and 2,250 trajectory points in 0.25 Hz. In Environment B, the exploration was completed in approximately 4 min, recording 212 points at 1 Hz.
Environment Traversing. Using the greedy TSP from Sec. III-B1, traversal trajectories were generated from the recorded exploration poses: 130 poses for Environment A and 15 poses for Environment B. Revisiting was significantly faster, taking approximately 35 min for Environment A and 2 min for Environment B, demonstrating the efficiency of the optimized traversal.
IV-B Real-time Semantic Mapping
For real-time semantic mapping, we equip Fetch with a laptop with an RTX 4090 GPU running Ubuntu 20.04 and ROS Noetic. The system then begins with the built occupancy map. After localizing using AMCL ROS [42], the robot progressively builds the semantic map. We set the object association distance threshold of 0.7 m for all categories. Subsequently, two additional iterations are performed, during which the environment is modified through the removal, addition, and/or relocating objects in the environments. Fig. 6 shows some of the changes in Environment B, reflected during semantic updating.
Env Class Iteration I (Initial run) Iteration II (vs I) Iteration III (vs I)
   GT Map $\DeltaGT Δ\DeltaMap$ $\DeltaGT Δ\DeltaMap$
A Table 45 45 0 -6 0* -8
Chair 109 84 -10 -11 0* -15
Door 170 153 0 +2 0 +3
B Person 2 2 0* 0 0 0
Trash Bin 2 2 -2 -2 -2 -1
Umbrella 1 1 -1 -1 -1 -1
Chair 3 3 0 0 -2 -2
Television 1 2 0 -1 0 -1
Cabinet 3 2 0 -1 0 0
[…]
@@ B Person 2 2 0* 0 0 0
Figure 6: Visualization of environmental changes in Environment B: Iteration II (2 trash bins removed) and Iteration III (2 chairs removed). When robot revisited, these changes are reflected in the updated semantic map.
IV-C Quantitative Analysis.
Table III reports the semantic mapping results across three iterations, with Iteration I serving as the reference. For Iteration I, we show the absolute ground truth (GT) objects and those mapped by the robot (Map). For Iterations II and III, we report only the relative changes ($\DeltaGT, Δ\DeltaMap$) compared to Iteration I.
[…]
Overall, this analysis shows that while stable objects (e.g., tables, doors) maintain consistent mapping across iterations, dynamic or occlusion-prone categories such as chairs exhibit greater variability in detection and mapping performance.
Class Iteration I Iteration II Iteration III
TP FP FN F1 TP FP FN F1 TP FP FN F1
Environment A
Table 36 9 9 0.8 36 3 9 0.86 33 4 12 0.80
Chair 84 0 25 0.87 72 1 27 0.84 67 2 42 0.75
Door 140 13 30 0.87 140 15 30 0.86 142 14 28 0.87
Environment B
Person 2 0 0 1.00 2 0 0 1.00 2 0 0 1.00
Trash Bin 2 0 0 1.00 0 0 0 - 0 1 0 0.00
Umbrella 1 0 0 1.00 0 0 0 - 0 0 0 -
Chair 3 0 0 1.00 3 0 0 1.00 1 0 0 1.00
Television 1 1 0 0.67 1 0 0 1.00 1 0 0 1.00
Cabinet 2 0 1 0.80 1 0 2 0.50 2 0 1 0.80
TABLE IV: Semantic map object count across Iteration I, II & III for A & B Environments. TP: True Positives, FP: False Positives, FN: False Negatives, F1 scores are reported.
Figure 7: Intermediate semantic maps generated by : (a) our method, (b) Khronos [26], running at 21FPS of data replay.
We compare and evaluate our semantic mapping approach with Khronos [26] on the real-world mezzanine scene dataset [43] provided as a ROS bag. Fig. 7 represents the semantic maps generated by both methods at an intermediate stage. The results are summarized in Table V. At 30 FPS of ROS bag replay, our method successfully maps 79 objects, while Khronos only maps 3, showing that our approach can process visual information faster and efficient as well. When the input rate is reduced to 21 FPS, Khronos maps 87 objects, while our method achieves a comparable 86.
Data Replay Rate Objects Mapped by Method
Khronos [26] Ours
21 FPS 87 86
30 FPS 3 79
TABLE V: Number of objects mapped by each method at different data FPS on the real-world mezzanine dataset.
This performance gap arises from the fundamental differences in system design. Since Khronos performs SLAM and semantic mapping in a single stage, its runtime is to an extent governed by loop-closure optimization, achieving superior SLAM performance, but limits the scalability of semantic mapping at higher frame rates. In contrast, our method decouples occupancy and semantic mapping, allowing it to achieve both higher throughput and competitive accuracy across different operating conditions.
IV-D Qualitative Analysis
@@
Fig. 8 illustrates several examples of accurate object detection during our experiments for chairs, tables, and doors.
@@
Limitations in Constrained Spaces. Narrow corridors limit the robot’s view of objects like doors, as seen in Fig. 9(f). Several of such doors are not detected, leading to a reduction in the number of object nodes in the semantic map. This is evident in Table III.
@@ V Conclusion and Future Work
We introduce a robotic system that enables a mobile robot to autonomously explore an unknown environment, recognize objects in the environment, and subsequently build a semantic map of the environment. The system utilizes a hybrid representation of the semantic map that consists of an occupancy grid map for geometry and a topological map for semantics. This representation enables us to update the semantic map efficiently to reflect the changes in the environment. Experiments conducted on a mobile manipulatory system in large- and medium-scale indoor environments demonstrate the effectiveness of our system in autonomous exploration and semantic mapping.
@@ Acknowledgment
This work was supported by the DARPA Perceptually-enabled Task Guidance (PTG) Program under contract number HR00112220005, the Sony Research Award Program, the National Science Foundation (NSF) under Grant Nos. 2346528 and 2520553, and the NVIDIA Academic Grant Program Award. The work of T. Summers was supported by the United States Air Force Office of Scientific Research under Grant FA9550-23-1-0424 and the National Science Foundation under Grant ECCS-2047040. We would like to thank our colleague, Jishnu Jaykumar P, for his assistance during the experiments.